# On the Alexander polynomial of special alternating links

### v2

@@ 3. $P_{D}(t)$ in terms of volumes of Root Polytopes of oriented co-Eulerian matroids
In this section, we prove Theorem 1.7. Before we do so, we review the work of Tóthmérész [21] on root polytopes of co-Eulerian matroids. We follow parts of her exposition.
A matroid $M$ is said to be regular if it is representable by a totally unimodular matrix, i.e. a matrix whose subdeterminants are all either $1$, 00, or $-1$. Let $A$ denote such a matrix and $\{a_{1},\dots,a_{m}\}$ its columns. For each circuit $C=\{i_{1},\dots,i_{j}\}$ of a regular matroid with a corresponding linear dependence relation $\sum_{k=1}^{j}\lambda_{k}a_{i_{k}}=0$ of columns of $A$, we may partition its elements into two sets: $C^{+}=\{i_{k}\ |\ \lambda_{k}>0\}$ and $C^{-}=\{i_{k}\ |\ \lambda_{k}<0\}$. Observe that scaling the expression $\sum_{k=1}^{j}\lambda_{k}a_{i_{k}}$ by a nonzero constant potentially interchanges $C^{+}$ and $C^{-}$, but the partition of $C$ into these two sets is well-defined up to this exchange. In this way, regular matroids are oriented.
@@
Tóthmérész demonstrates that for any pair of totally unimodular matrices $A$ and $A^{\prime}$ representing the same regular oriented matroid $M$, there is a lattice point-preseving linear bijection between $t\cdot\mathcal{Q}_{A}$ and $t\cdot\mathcal{Q}_{A^{\prime}}$ for any $t\in\mathbb{Z}$. In this way, the definition of a root polytope can be extended to regular oriented matroids. The set of arborescences of an Eulerian digraph yields a triangulation of the root polytope of the dual matroid. Recall that an arborescence of $D$ rooted at $r\in\operatorname{V}(D)$ is a directed subgraph $A$ such that for each other vertex $v$ of $D$, there is a unique directed path in $A$ from $r$ to $v$.
@@
Let $D$ be an Eulerian digraph, and let $A$ be any totally unimodular matrix representing the oriented dual of the oriented graphic matroid of $D$. Let $r\in V(D)$ and $\mathcal{D}=\{B\subset E(D)\ |\ E(D)-B\ \text{is an arborescence of}\ D\ \text{rooted at}\ r\}$. Then $\{\Delta_{B}\ |\ B\in\mathcal{D}\}$ is a triangulation of $\mathcal{Q}_{A}$.
@@
Let $D$ be an Eulerian digraph, and let $A$ be any totally unimodular matrix representing the oriented dual of the oriented graphic matroid of $D$. Let $r\in V(D)$. Then
@@
$c_{k}(D,r)=\sum_{i=0}^{k}(-1)^{i}\sum_{\begin{subarray}{c}\text{acyclic }E^{\prime}\subset E(D)\\ |E^{\prime}|=k-i\end{subarray}}\binom{|V(D)|-1-(k-i)}{i}\emph{Vol}(\mathcal{Q}_{H_{D/E^{\prime}}}).$
@@
Note that if $k>|V(D)|-1$, then $c_{k}(D,r)=0$ since a spanning tree in $D$ has $|V(D)|-1$ edges. Since #$\#col(A)-m=|V(D)|-1$ (both sides compute the size of a basis of $M$), the right-hand side also has no terms of degree greater than $|V(D)|-1$.
@@
for all $k\leq|V(D)|-1$. Since $c_{k}(D,r)=c_{|V(D)|-1-k}(D,r)$, this is also equivalent to
$\begin{split}c_{k}(D,r)&=\sum_{i=0}^{\#col(A)-|V(D)|+k-m+1}\sum_{\begin{subarray}{c}H\text{ with property *}\\ \#col(H)=|V(D)|-1-k+m+i\end{subarray}}(-1)^{i}\binom{|V(D)|-1-k+i}{i}Vol(Q_{H})\\ &=\sum_{i=0}^{k}\sum_{\begin{subarray}{c}H\text{ with property *}\\ \#col(H)=\#col(A)-k+i\end{subarray}}(-1)^{i}\binom{|V(D)|-1-k+i}{i}Vol(Q_{H})\end{split}$
Recall that by standard results in matroid theory, deleting elements of $M^{*}$ (or columns of $A$) is equivalent to contracting the corresponding elements in $M$. Furthermore, the set of columns we remove from $A$ to form $H$ are disjoint to a basis of $M^{*}$ (since removing them did not decrease the rank of the matrix), so these columns form an independent set of $M$. Since the independent sets in the graphic matroid of $D$ are precisely acyclic subsets of edges of $D$, the desired result is equivalent to
$c_{k}(D,r)=\sum_{i=0}^{k}(-1)^{i}\sum_{\begin{subarray}{c}\text{acyclic }E^{\prime}\subset E(D)\\ |E^{\prime}|=k-i\end{subarray}}\binom{|V(D)|-1-(k-i)}{i}\emph{Vol}(\mathcal{Q}_{H_{D/E^{\prime}}})$
@@
Both Theorem 1.7 and the following corollary are closely related to the beautiful unpublished work of Li and Postnikov [12].
@@
(5) $\Delta_{L_{G}}(-t)=\sum_{\begin{subarray}{c}H\subset G\\ H\text{ connected}\end{subarray}}Vol(Q_{H})(q-1)^{|E(H)|-|V(G)|+1}.$
@@
(6) $P_{D}(t)=\sum_{\begin{subarray}{c}H\subset G\\ H\text{ connected}\end{subarray}}Vol(Q_{H})(q-1)^{|E(H)|-|V(G)|+1}.$
@@ 4. $P_{D}(t)$ as the Alexander polynomial of special alternating links
[16, Theorem 2] For an alternating dimap $D$, the Alexander polynomial of $D$, $P_{D}(t)$, equals the Alexander polynomial $\Delta_{L_{G}}(-t)$ for the special alternating link $L_{G}$ associated to the planar dual $G$ of $D$. In other words:
@@
Theorem 1.4 is stated in [16] as $\Delta_{L_{G}}(-t)\sim P_{D}(t)$. By Theorem 1.3, $P_{D}(t)$ is palindromic; therefore, $\Delta_{L_{G}}(-t)\sim P_{D}(t)$ is equivalent to $\Delta_{L_{G}}(-t)\sim\sum_{k=0}^{|\operatorname{V}(D)|-1}c_{k}(D,r)t^{|\operatorname{V}(D)|-1-k}$ which is what we state in Theorem 1.4.
@@ 4.1. Kauffman’s state model
The State Summation, introduced by Kauffman [11], describes the Alexander polynomial as a sum over weights of decorations of a link diagram, called states [11]. To form a state, one begins with a (not necessarily alternating) link diagram and marks two (arbitrary) adjacent regions. Each meeting between a nonmarked region and a crossing is labeled as follows:
@@
Each state is then a bijection between crossings in the diagram and the nonmarked regions so that each crossing is mapped to one of the four regions it meets. This bijection is presented diagrammatically, as in Figure 2, via state markers. For each state $S$, there is a signed monomial $\langle L|S\rangle$, computed as the product of the weights corresponding to the marked region at each crossing.
@@
Figure 2. The four states of the positive $K(2,2)$ along with their monomial weights and the resulting Alexander polynomial.
@@
$\Delta_{L}(t)\sim\sum_{S}\langle L|S\rangle(-1)^{b(S)}$.
@@
Kauffman provides a bijection between states of a link diagram and spanning trees of the checkerboard graph (or equivalently, spanning trees of the checkerboard dual). Given a state $S$, one draws an edge between vertices in its associated tree if and only if they are separated by a crossing whose state marker points from the region corresponding to one vertex to the region corresponding to the other. In this way, a state identifies a pair of spanning trees that we refer to as dual spanning trees: one in $G$ and one in $G^{*}$. We note that these dual spanning trees are not planar duals of each other; rather, they correspond to bases in the graphic matroids of $G$ and $G^{*}$ which are complements of one another.
Conversely, given a spanning tree $T$ of $G$ and its dual spanning tree $T^{\prime}$ in $G^{*}$ (which consists of the edges in $E(G^{*})$ which do not correspond to edges in $E(T)$, see Figure 3), one recovers a state by assigning edge orientations that make $T$ and $T^{\prime}$ oriented spanning trees rooted at the vertices corresponding to the two adjacent marked regions. When a directed edge is incident to a crossing, place a state marker pointing into that crossing from the region corresponding to the initial vertex of the edge, as in Figure 3.
@@
We first consider the case where $G$ is a bipartite cycle. It is easy to see that the number of edges $(a,b)$ with $a\in A$ and $b\in B$ is $|A|$ if $r^{\prime}\in B$ and $|A|-1$ if $r^{\prime}\in A$. In either case, this is independent of the choice of spanning tree.
Next, let $G$ be an arbitrary bipartite graph, and let $r^{\prime}\in\operatorname{V}(G)$. Any two spanning trees of $G$ are connected by a sequence of edge exchanges. It, thus, suffices to show that if $S$ and $S^{\prime}$ are two spanning trees related by a single edge exchange and oriented so that they are oriented spanning trees rooted at $r^{\prime}$, then they have the same number of oriented edges $(a,b)$ with $a\in A$ and $b\in B$.
Notice that $S\cup S^{\prime}$ contains a single undirected cycle $C$ with distinct edges $e,e^{\prime}\in\operatorname{E}(C)$ such that $(S\cup S^{\prime})\setminus\{e\}=S$ and $(S\cup S^{\prime})\setminus\{e^{\prime}\}=S^{\prime}$. Consider the forest $(S\cup S^{\prime})\setminus C$. The component of this forest containing $r^{\prime}$ intersects $C$ in at least one vertex. Indeed, it must intersect $C$ in exactly one vertex $s$ – if it intersected $C$ in at least two distinct vertices, then $S\cup S^{\prime}$ would admit two distinct (undirected) cycles. Treating $C\setminus\{e\}$ and $C\setminus\{e^{\prime}\}$ as oriented spanning trees rooted at $s$, the above argument shows that they have the same number of oriented edges $(a,b)$ where $a\in A$ and $b\in B$. Since $S\setminus C=S^{\prime}\setminus C$, we conclude that $S$ and $S^{\prime}$ have the same number of oriented edges $(a,b)$ where $a\in A$ and $b\in B$. ∎
@@
[…]
@@ 5. A note on generalizing Fox’s conjecture for special alternating links
The Laplacian matrix of an Eulerian graph $D$ on the vertex set $[n]=\{1,\ldots,n\}$ is the matrix $L(D)=(l_{i,j})_{i,j\in[n]}$, defined by
$l_{i,j}=\begin{cases}\operatorname{outdeg}(i)&\text{for}\ i=j\\ |\{e\in E(D)\ :\ $e$\ \text{is an edge from}\ i\ \text{to}\ j\}|&\text{for}\ i\neq j\end{cases}.$
Let $\overline{L(D)}$ be the reduced Laplacian obtained from $L$ by removing a row and a column (of the same index).
@@
Figure 7. The oriented three-cycle $D$, superimposed on a link diagram of the trefoil knot, and its checkerboard graph.
@@ 6. Bijection between the Crowell and Kauffman state models
This section is devoted to relating the first state model for the Alexander polynomial: Crowell’s model [3] to the most well-known one: the Kauffman state model [11]. In Theorem 6.5, we define a weight-preserving bijection between the states of the Crowell model and the states of the Kauffman model for special alternating links.
We will use the following notation throughout Section 6. We denote by $G$ the planar bipartite graph which is the checkerboard graph of the special alternating link $L:=L_{G}$. The alternating dimap $D$ is the directed Eulerian dual of $G$ (see Section 1.2 and Figure 1).

### v3

@@ 3. $P_{D}(t)$ in terms of volumes of Root Polytopes of oriented co-Eulerian matroids
In this section, we prove Theorem 1.7. Before we do so, we review the work of Tóthmérész [23] on root polytopes of co-Eulerian matroids, following parts of her exposition.
A matroid $M$ is said to be regular if it is representable by a totally unimodular matrix, i.e. a matrix whose subdeterminants are all either $1$, 00, or $-1$. Let $A$ denote such a matrix and $\{a_{1},\dots,a_{m}\}$ its columns. For each circuit $C=\{i_{1},\dots,i_{j}\}$ of a regular matroid with a corresponding linear dependence relation $\sum_{k=1}^{j}\lambda_{k}a_{i_{k}}=0$ of columns of $A$, we may partition its elements into two sets: $C^{+}=\{i_{k}\ |\ \lambda_{k}>0\}$ and $C^{-}=\{i_{k}\ |\ \lambda_{k}<0\}$. Observe that scaling the expression $\sum_{k=1}^{j}\lambda_{k}a_{i_{k}}$ by a nonzero constant potentially interchanges $C^{+}$ and $C^{-}$, but the partition of $C$ into these two sets is well-defined up to this exchange. In this way, regular matroids can be oriented.
@@
Tóthmérész demonstrates that for any pair of totally unimodular matrices $A$ and $A^{\prime}$ representing the same regular oriented matroid $M$, there is a lattice point-preseving linear bijection between $t\cdot\mathcal{Q}_{A}$ and $t\cdot\mathcal{Q}_{A^{\prime}}$ for any $t\in\mathbb{Z}$. The definition of a root polytope can thus be extended to regular oriented matroids by considering the root polytope of a totally unimodular matrix representing the matroid. The set of arborescences of an Eulerian digraph yields a triangulation of the root polytope of the dual matroid. Recall that an arborescence of $D$ rooted at $r\in\operatorname{V}(D)$ is a directed subgraph $A$ such that for each other vertex $v$ of $D$, there is a unique directed path in $A$ from $r$ to $v$.
@@
Let $D$ be an Eulerian digraph, and let $A$ be any totally unimodular matrix representing the oriented dual of the oriented graphic matroid of $D$. Let $r\in\operatorname{V}(D)$ and $\mathcal{D}=\{B\subset\operatorname{E}(D)\ |\ \operatorname{E}(D)-B\ \text{is an arborescence of}\ D\ \text{rooted at}\ r\}$. Then $\{\Delta_{B}\ |\ B\in\mathcal{D}\}$ is a triangulation of $\mathcal{Q}_{A}$.
@@
Let $D$ be an Eulerian digraph, and let $A$ be any totally unimodular matrix representing the oriented dual of the oriented graphic matroid of $D$. Let $r\in\operatorname{V}(D)$. Then
@@
$c_{k}(D,r)=\sum_{i=0}^{k}(-1)^{i}\sum_{\begin{subarray}{c}\text{cycle-free }E^{\prime}\subset\operatorname{E}(D)\\ |E^{\prime}|=k-i\end{subarray}}\binom{|\operatorname{V}(D)|-1-(k-i)}{i}\emph{Vol}(\mathcal{Q}_{H_{D/E^{\prime}}}).$
@@
Note that if $k>|\operatorname{V}(D)|-1$, then $c_{k}(D,r)=0$ since a spanning tree in $D$ has $|\operatorname{V}(D)|-1$ edges. Since #$\#col(A)-m=|\operatorname{V}(D)|-1$ (both sides compute the size of a basis of $M$), the right-hand side also has no terms of degree greater than $|\operatorname{V}(D)|-1$.
@@
for all $k\leq|\operatorname{V}(D)|-1$. Since $c_{k}(D,r)=c_{|\operatorname{V}(D)|-1-k}(D,r)$, this is also equivalent to
$\begin{split}c_{k}(D,r)&=\sum_{i=0}^{\#col(A)-|\operatorname{V}(D)|+k-m+1}\sum_{\begin{subarray}{c}H\text{ with property *}\\ \#col(H)=|\operatorname{V}(D)|-1-k+m+i\end{subarray}}(-1)^{i}\binom{|\operatorname{V}(D)|-1-k+i}{i}Vol(Q_{H})\\ &=\sum_{i=0}^{k}\sum_{\begin{subarray}{c}H\text{ with property *}\\ \#col(H)=\#col(A)-k+i\end{subarray}}(-1)^{i}\binom{|\operatorname{V}(D)|-1-k+i}{i}Vol(Q_{H})\end{split}$
Recall that by standard results in matroid theory, deleting elements of $M^{*}$ (or columns of $A$) is equivalent to contracting the corresponding elements in $M$. Furthermore, the set of columns we remove from $A$ to form $H$ are disjoint to a basis of $M^{*}$ (since removing them did not decrease the rank of the matrix), so these columns form an independent set of $M$. Since the independent sets in the graphic matroid of $D$ are precisely cycle-free subsets of edges of $D$, the desired result is equivalent to
$c_{k}(D,r)=\sum_{i=0}^{k}(-1)^{i}\sum_{\begin{subarray}{c}\text{cycle-free }E^{\prime}\subset\operatorname{E}(D)\\ |E^{\prime}|=k-i\end{subarray}}\binom{|\operatorname{V}(D)|-1-(k-i)}{i}\emph{Vol}(\mathcal{Q}_{H_{D/E^{\prime}}})$
@@
(7) $\Delta_{L_{G}}(-t)\sim\sum_{\begin{subarray}{c}H\subset G\\ H\text{ connected}\end{subarray}}Vol(Q_{H})(t-1)^{|\operatorname{E}(H)|-|\operatorname{V}(G)|+1}.$
@@
(8) $P_{D}(t)=\sum_{\begin{subarray}{c}H\subset G\\ H\text{ connected}\end{subarray}}Vol(Q_{H})(t-1)^{|\operatorname{E}(H)|-|\operatorname{V}(G)|+1}.$
@@
As explained in Section 1.5, Theorem 1.7 was inspired by the work of Li and Postnikov [14], and Corollary 1.8 can also be deduced from results of Li and Postnikov [14] and Kálmán, Mészáros, Postnikov [10].
@@ 4. $P_{D}(t)$ as the Alexander polynomial of special alternating links
[18, Theorem 2] For an alternating dimap $D$, the Alexander polynomial of $D$, $P_{D}(t)$, equals the Alexander polynomial $\Delta_{L_{G}}(-t)$ for the special alternating link $L_{G}$ associated to the planar dual $G$ of $D$. In other words:
@@
Theorem 1.4 is stated in [18] as $\Delta_{L_{G}}(-t)\sim P_{D}(t)$. By Theorem 1.3, $P_{D}(t)$ is palindromic; therefore, $\Delta_{L_{G}}(-t)\sim P_{D}(t)$ is equivalent to $\Delta_{L_{G}}(-t)\sim\sum_{k=0}^{|\operatorname{V}(D)|-1}c_{k}(D,r)t^{|\operatorname{V}(D)|-1-k}$, the expression in Theorem 1.4.
@@ 4.1. Kauffman’s state model
The State Summation, introduced by Kauffman [13], describes the Alexander polynomial as a sum over weights of decorations of a link diagram, called states [13]. We begin with a (not necessarily alternating) link diagram $L$ and marks two (arbitrary) adjacent regions. At every crossing, we assign a weight to each nonmarked region that meets it, according to the following convention:
@@
A state is a bijection between crossings in the diagram and the nonmarked regions so that each crossing is mapped to one of the four regions it meets. This bijection is presented diagrammatically, as in Figure 2, by drawing a state marker at each crossing to show which region it is mapped to. For each state $S$, there is a signed monomial $\langle L|S\rangle$, computed as the product of the weights corresponding to the region with the state marker at each crossing.
@@
Figure 2. The four states of the positive $K(2,2)$ along with their monomial weights and the resulting Alexander polynomial. The two adjacent marked regions are each labeled with an x.
@@
For any link $L$, $\Delta_{L}(t)\sim\sum_{S}\langle L|S\rangle(-1)^{b(S)}$.
@@
Kauffman provides a bijection between states of a link diagram and spanning trees of the checkerboard graph (or equivalently, spanning trees of the checkerboard dual). Given a state $S$, one draws an edge between a pair of vertices in its associated tree if and only if the regions corresponding to those vertices are separated by a crossing whose state marker is in one of those two regions (see Figure 3). In this way, a state identifies a pair of spanning trees that we refer to as dual spanning trees: one in $G$ and one in $G^{*}$. We note that these dual spanning trees are not planar duals of each other; rather, they correspond to bases in the graphic matroids of $G$ and $G^{*}$ that are complements of one another.
Conversely, given a spanning tree $T$ of $G$ and its dual spanning tree $T^{\prime}$ in $G^{*}$ (which consists of the edges in $\operatorname{E}(G^{*})$ which do not correspond to edges in $\operatorname{E}(T)$, see Figure 3), one recovers a state by assigning edge orientations that make $T$ and $T^{\prime}$ oriented spanning trees rooted at the vertices corresponding to the two adjacent marked regions. When a directed edge is incident to a crossing, place a state marker pointing into that crossing from the region corresponding to the initial vertex of the edge, as in Figure 3.
@@
Since the color classes $A$ and $B$ form a proper 2-coloring of $G$, every edge with initial vertex in $A$ has the form $(a,b)\in\operatorname{E}(S)$ where $a\in A$ and $b\in B$. If $S$ is an oriented spanning tree, then by definition, every non-root vertex $v\in\operatorname{V}(G)$ is the initial vertex of exactly one edge of $S$. In particular, the number of edges $(a,b)$ with $a\in A$ and $b\in B$ is precisely the number of non-root vertices in $A$. This number is $|A|$ if $r^{\prime}\in B$ and $|A|-1$ if $r^{\prime}\in A$. In either case, this is independent of the choice of spanning tree.
@@
The proof of Theorem 1.4 shows that in the case of special alternating links $L_{G}$, the Alexander polynomial $\Delta_{L_{G}}(-t)=P_{D}(t)$ can be calculated by collecting terms $t^{|\operatorname{E}(T)\setminus\operatorname{E}_{\text{away}}(T)|}$ for each spanning tree $T$ of $D$ (with specified arbitrary root $r$). We will later use the notation $\operatorname{wt}_{K}(T)=t^{|\operatorname{E}(T)\setminus\operatorname{E}_{\text{away}}(T)|}$ where the subscript $K$ indicates that we are dealing with the Kauffman weight for special alternating links.
@@ 5. A note on generalizing Fox’s conjecture for special alternating links
The Laplacian matrix of an Eulerian digraph $D$ on the vertex set $[n]=\{1,\ldots,n\}$ is the matrix $L(D)=(l_{i,j})_{i,j\in[n]}$, defined by
$l_{i,j}=\begin{cases}\operatorname{outdeg}(i)&\text{for}\ i=j\\ |\{e\in\operatorname{E}(D)\ :\ $e$\ \text{is an edge from}\ i\ \text{to}\ j\}|&\text{for}\ i\neq j\end{cases}.$
Let $\overline{L(D)}$ be the reduced Laplacian obtained from $L$ by removing a row and a column (of the same index). We will make use of the following straightforward consequence of the matrix-tree theorem.
@@
Figure 7. The oriented three-cycle $D$, superimposed on a link diagram of the trefoil knot and its checkerboard graph.
@@ 6. Bijection between the Crowell and Kauffman state models
This section is devoted to relating the first state model discovered for the Alexander polynomial: Crowell’s model [4] to the most well-known one: the Kauffman state model [13]. In Theorem 6.7, we define a weight-preserving bijection between the states of the Crowell model and the states of the Kauffman model for special alternating links.
Throughout this section, $G$ denotes the planar bipartite graph which is the checkerboard graph of the special alternating link $L:=L_{G}$, and $D$ denotes the directed Eulerian dual of $G$ (see Section 1.2 and Figure 1).